OpenStudy (anonymous):

solve the system of equations for x : 2x+y=4 x-2y=7

OpenStudy (chris):

what do you get when you multiply the top equation by 2?

OpenStudy (chris):

i.e. 2(2x + 4) = 2*4, so 4x+2y=8

OpenStudy (anonymous):

so whats the answer to x ?

OpenStudy (chris):

Now given: 4x + 2y = 8, "add" x - 2y = 7. and you end up canceling out the y's. So 5x = 15. And x = 3